The sum of two numbers is -12.43 and their difference is 117.77

(I tried adding two negative numbers to equal -12.43, but nothing large enough to have a difference of 117.77.

Answer by mananth(16949)

x+y = -12.43
...
x-y = 117.77
..
add the two equations
2x=105.34
/2
x=52.67
...
plug value of x in first equation
52.67+y=-12.43
-52.67
y=-12.43-52.67
y = -65.10
...
m.ananth@hotmail.ca